OpenStudy (anonymous):

HELP PLEASE Given the indicated parts of triangle ABC with γ = 90°, approximate the remaining parts. (Round your answers for b and c to one decimal place. β = 68°50', a = 20.1

OpenStudy (anonymous):

Can you please draw a diagram

OpenStudy (anonymous):

maybe its like thisuse trigonometric relationship in a right triangle\[\sin \beta = \frac{b}{c} \]\[\tan \beta = \frac{b}{a}\]\[\cos \beta= \frac{a}{c}\]